# Through the Judge’s Eyes: Inferred Thinking Traces Improve Reliability of LLM Raters

### v1

@@ Appendix B SFT Details
Appendix C Prompts
Instruction refinement ¡instruction¿ You will be given a list of analysis on scoring the {dimension} of {generation_type}. Your task is to extract a concrete and concise step-by-step instruction from the analysis that could be easily followed by an annotator without any training. Based on your extraction, improve the original annotation guidelines (¡original_codebook¿…¡/original_codebook¿). Only improve the instruction part (e.g. thinking path to follow), do not change the rubric part (e.g. criteria). Your final annotation guidelines should be put into ¡codebook¿…¡/codebook¿ tag. ¡/instruction¿ ¡original_codebook¿ {original_codebook} ¡/original_codebook¿ ¡analysis¿ {analysis} ¡/analysis¿
Critique extraction from CoT ¡instruction¿ You will be given an analysis on scoring the {dimension} of a {generation_type}. Your task is to list all the critique from the analysis that help evaluate the {dimension} of the {generation_type}. Your response should be a list of sentences, each on a new line, without bullet points. Your response should not include any other text. ¡/instruction¿ ¡analysis¿ {cot} ¡/analysis¿
Rubric refinement ¡instruction¿ You will be shown a list of diverse critiques on the {dimension} of {generation_type}. Each of the critique is corresponding to a specific rating level. Your task is to summarize the critiques into a single rubric, based on the pattern of the different rating levels. Based on the new criteria, improve the original annotation guidelines. Only modify the criterion part. Your final annotation guidelines should be put into ¡codebook¿…¡/codebook¿ tag. ¡/instruction¿ ¡original_codebook¿ {original_codebook} ¡/original_codebook¿ ¡critiques¿ {critiques} ¡/critiques¿
Appendix D Complete Examples of Annotation Codebooks
A complete list of codebooks before and after refinement is provided in our codebase: https://github.com/xingjian-zhang/thru_judge_eye. Here we show a complete refined version of Figure 4.
@@ Appendix D Complete Examples of Annotation Codebooks
(1) The story is very simple, lacks depth, and introduces no intricate or developed elements. (2) The story is somewhat simple, with few elements that are underdeveloped, fragmented, or disconnected. (3) The story demonstrates some complexity, introducing multiple elements, but lacks depth, development, or coherence in integrating these elements. (4) The story is mostly complex, weaving together several developed elements with minor areas that could be expanded or refined. (5) The story is highly elaborate and complex, integrating multiple layers (e.g., detailed world-building, character depth, advanced structure, thematic sophistication) in a cohesive and interconnected way.

### v2

@@ Appendix C SFT Details
Appendix D Prompts
Instruction refinement <instruction> You will be given a list of analysis on scoring the {dimension} of {generation_type}. Your task is to extract a concrete and concise step-by-step instruction from the analysis that could be easily followed by an annotator without any training. Based on your extraction, improve the original annotation guidelines (<original_codebook>…</original_codebook>). Only improve the instruction part (e.g. thinking path to follow), do not change the rubric part (e.g. criteria). Your final annotation guidelines should be put into <codebook>…</codebook> tag. </instruction> <original_codebook> {original_codebook} </original_codebook> <analysis> {analysis} </analysis>
Critique extraction from CoT <instruction> You will be given an analysis on scoring the {dimension} of a {generation_type}. Your task is to list all the critique from the analysis that help evaluate the {dimension} of the {generation_type}. Your response should be a list of sentences, each on a new line, without bullet points. Your response should not include any other text. </instruction> <analysis> {cot} </analysis>
Rubric refinement <instruction> You will be shown a list of diverse critiques on the {dimension} of {generation_type}. Each of the critique is corresponding to a specific rating level. Your task is to summarize the critiques into a single rubric, based on the pattern of the different rating levels. Based on the new criteria, improve the original annotation guidelines. Only modify the criterion part. Your final annotation guidelines should be put into <codebook>…</codebook> tag. </instruction> <original_codebook> {original_codebook} </original_codebook> <critiques> {critiques} </critiques>
Appendix E Complete Examples of Annotation Codebooks
A complete list of codebooks before and after refinement is provided in our anonymous codebase: https://anonymous.4open.science/r/thru_judge_eye-56DD/codebooks/. Here we show a complete refined version of Figure 4.
@@ Appendix E Complete Examples of Annotation Codebooks
1. The story is very simple, lacks depth, and introduces no intricate or developed elements. 2. The story is somewhat simple, with few elements that are underdeveloped, fragmented, or disconnected. 3. The story demonstrates some complexity, introducing multiple elements, but lacks depth, development, or coherence in integrating these elements. 4. The story is mostly complex, weaving together several developed elements with minor areas that could be expanded or refined. 5. The story is highly elaborate and complex, integrating multiple layers (e.g., detailed world-building, character depth, advanced structure, thematic sophistication) in a cohesive and interconnected way.